OpenStudy (anonymous):

(MEDALS!!!!!!!!) Please answer. Solve. 6/11x + 4 = 2/11x + 16

OpenStudy (hitaro9):

6/11x +4 = 2/11x+ 16 subtracting 4 from both sides 6/11 x +4 -4 = 2/11 x +16 -4 6/11x = 2/11x +12 Multiple both sides by 11/6 (11/6)(6/11)x = (11/6)(2/11x+12) x= (1/3)x + 12(11/6) x = (1/3)x + 22 subtract 1/3 x from both sides 2/3(x) = 22 Multiply both sides by 3/2 x = 33

OpenStudy (hitaro9):

You can plug that in to both sides and end up getting the same answer on both sides, so its valid.